Q2.Calculate Elasticity under following questions:
a) Price of Pepsi is Rs.120 and its quantity demanded is 300 bottles. Suppose due to Raman Pepsi
company reduces price of it to Rs. 108 ,due to it its demand increases to 360 bottles. Calculate
elasticity and also mention the type of elasticity.
b) Mr.Ali has an income of Rs.30,000 and he purchases 100 units of X-GOOD. His income decreases
by 20% and now he can purchase 90 units of X-GOOD.Calculate Calculate elasticity and also
mention the type of elasticity.
c) Calculate elasticity and also mention the type of elasticity of the following:

Price of Milk Pack (1litre) Quantity demanded of Olpers Milk
Rs.140 200 liters
120 160
1
Expert's answer
2020-04-29T09:22:36-0400


a) Price of Pepsi is Rs.120 and its quantity demanded is 300 bottles. Suppose due to Raman Pepsi

company reduces price of it to Rs. 108 ,due to it its demand increases to 360 bottles. Calculate

elasticity and also mention the type of elasticity.


The price elasticity of demand will e given by:



"E_d = \dfrac{\Delta Q}{\Delta P}\cdot \dfrac{(P_1+ P_2)/2}{(Q_1+Q_2)/2}""E_d = \dfrac{360 - 300}{108 - 120}\cdot \dfrac{(108+120)/2}{(360+300)/2}"

"E_d = \dfrac{60}{-12}\cdot \dfrac{114}{330}"

"E_d \approx -1.73"

"|E_d| \approx 1.73"

The demand is elastic since the elasticity is greater than 1.



b) Mr.Ali has an income of Rs.30,000 and he purchases 100 units of X-GOOD. His income decreases by 20% and now he can purchase 90 units of X-GOOD. Calculate elasticity and also

mention the type of elasticity.


The income elasticity of demand is given by:



"E_Y = \dfrac{\%\Delta Q}{\%\Delta Y}"

The percentage change in the demand for good X is:


"\%\Delta Q = \dfrac{90 - 100}{100}\times 100 = -10\%"

If the change in income "\%\Delta Y = 20\%" , then the income elasticity of demand is:



"E_Y = \dfrac{-10\%}{-\%20} = 0.5"

c) Calculate elasticity and also mention the type of elasticity of the following:

Price of Milk Pack (1litre) Quantity demanded of Olpers Milk

Rs.140 160liters

Rs. 120 200 liters


The price elasticity of demand for milk will be given by:



"E_d = \dfrac{\Delta Q}{\Delta P}\cdot \dfrac{(P_1+ P_2)/2}{(Q_1+Q_2)/2}"

"E_d = \dfrac{200 - 160}{120 - 140}\cdot \dfrac{(120+140\ )/2}{(160+200)/2}"

"E_d = \dfrac{40}{-20}\cdot \dfrac{130}{180}"

"E_d \approx -1.44""|E_d| \approx 1.44"


The demand is elastic since the elasticity is greater than 1.